the length of a rectangle is 5 feet longer than its width. If its area is 36 square feet, find its perimeter.
Answer by TimothyLamb(4379)

x = length (feet)
y = width (feet)
---
x = y + 5
y = x - 5
xy = 36
x(x - 5) = 36
xx - 5x - 36 = 0
---
the above quadratic equation is in standard form, with a=1, b=-5 and c=-36
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 -5 -36
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = 9
x = -4
---
the negative root doesn't fit the problem statement, so use the positive root:
---
x = 9
y = 9 - 5
y = 4
---
answer:
p = 2x + 2y
p = 2*9 + 2*4
p = 26 feet
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---